OpenStudy (anonymous):

The lengths of the sides of a right triangle are a and b, and the hypotenuse is c. Find the area of the triangle. a = 5 ft.; c = 13 ft. A = _____sq. ft.

OpenStudy (anonymous):

A squared + B squared = c squared

OpenStudy (anonymous):

sub the #s you have in for that

OpenStudy (anonymous):

here is the file

OpenStudy (anonymous):

Pythagorean Theorem c^2 = a^2 + b^2 13^2 = 5^2 + b^2 169 = 25 + b^2 144 = b^2 b = 12 This is also a pythagorean triple.

OpenStudy (ajprincess):